OpenStudy (anonymous):

Simplify showing your work. Remember to use a ^ symbol to indicate exponents. (3pr^3)^2(pr)=

OpenStudy (anonymous):

I know the answer is 9p^3r^7 but i dont know how to work it out

OpenStudy (mathstudent55):

You need these two rules: When you raise a product to a power, raise every factor to the power. \((ab)^m = a^mb^m\) When you raise a power to a power, multiply the powers. \( (a^m)^n = a^{mn} \)

OpenStudy (mathstudent55):

\( (3pr^3)^2(pr) \) \( = 3^2 p^2 (r^3)^2 (pr) \) \( = 9 p^2p r^6 r \) \( =9 p^3 r^7 \)